Completed-Planned maintenance:Computing Centre UPS systems-23 Aug

This work was successfully completed.

On Thursday 23 August between 10am and 4pm we are replacing the batteries in one of our two Uninterruptible Power Supply (UPS) systems that provide battery backup for our Core Equipment Room in Park Place.

During this time our services will continue to run as normal, but we are at a higher risk of loss of service should we experience a campus power failure that affects the Computing Centre. Estates & Buildings are aware of this work and have assured us that there is no other planned power work during this time.

If you have any questions or concerns please contact Service Desk quoting Change No 91.

Completed-Planned maintenance:Network system migration

We are planning to migrate departments to a new system that changes the way device configuration information is used by the network. This will provide a more resilient service and an efficient way of managing the space we have on our network for devices.  At the same time we will be enhancing security on this aspect of the service.

Our technical staff will be in touch with departmental IT support contacts to gather information and consult on an appropriate time to do this work.

No: 45830 – Issues with new passwords

We are aware that some users are experiencing issues accessing dMail once they have changed their password. This is due to users entering the same password as before into the new password field.

Although the Password Management Portal will technically accept the same password, the directory used to authorise access to dMail will not allow this to synchronise and so users are experiencing errors.

When changing your password please choose a new password from before.

Please contact Service Desk if you have any questions or concerns.

Resolved-No:45354-Issues with directory link to dMail

This issue has now been resolved and synchronisation between our directory and dMail has been restored.

***************

14.59 – 08/08/12

We are aware that there is an issue with the linkage between our University directory and dMail. This means that new user accounts are not being created and changes to existing staff or students’ accounts are not being updated in dMail.

We have reported this problem to our supplier and expect this to be resolved later this afternoon.

We apologise for any inconvenience this causes.

If you have any questions or concerns contact Service Desk quoting reference no 45354.

No: 45109 – Campus Power Outages Saturday 4th Aug – 07.30am

Information from Estates and Buildings – The power is to be switched off to a number of buildings on the Campus to undertake a high voltage safety inspection.  The outage will occur in the following buildings at 07.30am for approximately 15-20mins.

ISE (Institute for Sports and Exercise)

Belmont Tower

QMB (Queen Mother Building)

If you have any issues regarding this, please contact Estates and Buildings direct, or if any IT issues arise please contact Service Desk on Monday 6th Aug from 08.45am on ext 88000.

 

Change to VPN connections-Tuesday 21 August

On Tuesday 21 August 2012 we will be changing the way users connect to VPN.

VPN users will have to change their VPN settings in advance and we will be in touch with guidance on how to do this in w/b 6 August. We cannot provide the new settings until a network change is made next week, but we are keen that users have sufficient notice of the planned change to this service.

If you have any questions or concerns please contact Service Desk and quote reference 44588.

Network migration to JANET – 31 July 2012

UPDATE: 1st August 9.40 a.m.

We have now completed migration of our network infrastructure from FaTMAN to JANET UK.

If you have any questions about this change, please contact Service Desk,  quoting Change Ref 82.

**************************************************************

UPDATE: 31 July 11.30a.m.

An issue with the firewall was identified this morning. This has delayed progress with this work until Wednesday 1 August between 7.00 and 8.00 a.m.

**********************

On Tuesday 31 July 2012 users will be unable to connect to the network for a short period between 7.00 and 8.30, and the network should be considered at risk during this time.

During the at risk time slot we will migrate our network infrastructure from FaTMAN to JANET UK. The JANET network has the potential to offer internet connections of up to 10G.

If you have any questions or concerns please contact Service Desk and quote change ref 82.

Changes to Service Desk notifications-No:44742

Users will notice that the notifications that they receive from Service Desk will be different this morning. We have updated our notification emails to prepare for the rollout of our self-service portal, which be tried out by Local Analysts and Local Support Contacts very soon.

If you have any feedback, questions or concerns please contact Service Desk quoting reference 44742.